# NMM-HRI: Natural multimodal Human-Robot Interaction with Voice and Deictic Posture via Large Language Model

### v2

@@ I INTRODUCTION
With an aging population, the demand for efficient caregiving solutions is growing, yet labor costs continue to increase. Robotics manipulators have shown promise in alleviating these challenges by automating tasks that traditionally require human intervention [1]. However, one of the most critical and unresolved issues is ensuring effective Human-Robot Interaction (HRI) for elderly users. For HRI systems to truly benefit this population, they must offer interaction methods that are intuitive and natural, resembling everyday human communication. Current systems often rely on humans to memorize complex language syntax or master complex hand gestures [2], which are impractical for the elderly. This highlights the urgent need for a simpler yet highly effective method that allows robots to understand and execute commands from elderly users with ease and reliability.
In the past year, large language models (LLMs) have emerged as promising tools for HRI [3]. Their advanced reasoning and language capabilities make them promising for improving communication between humans and robots. However, directly applying LLMs to HRI presents several challenges. First, LLMs often require users to input detailed and structured text commands, which can be tedious and difficult to understand. Second, without integrated sensing capabilities, LLMs struggle to comprehend the environmental context or specific actions, limiting their effectiveness in real-world applications. Finally, LLMs are prone to hallucinations, generating inaccurate or unsafe responses, which can lead to harmful outcomes when used in control systems without close monitoring. These challenges highlight the need for a more robust integration of LLMs into HRI systems.
@@ II RELATED WORK
As perception and navigation solutions advance, robots are increasingly integrated into daily life. However, robots designed for complex tasks, such as surgical [4] and rehabilitation applications [5], face stringent safety requirements and must be validated for many years before widespread use in real world scenarios. In contrast, service robots, particularly for elderly and patient care, demonstrate greater immediate potential. However, current service robot systems often rely on single-modal inputs, such as pure hand gestures [2], voice commands [6], or combinations of body language such as hand gestures and body postures [7]. Hand gesture-based HRI methods [2], for example, frequently use the Leap Motion sensor to detect hand movements and translate them into commands. However, these sensors have a limited field of view, restricting their applicability to broader environments. To address this, some researchers [7] have introduced Kinect cameras to extend the range of the sensor and detect both human posture and gestures. Nonetheless, for elderly users, accurately memorizing and executing the required gestures for commands remains a significant challenge, limiting the effectiveness of these approaches. Voice commands have always been a natural candidate for HRI problems. People often employ various approaches, such as LSTM, hidden Markov models, or an attention-based encoder-decoder network [8], to parse voices into actionable commands. One good example is the Amazon Alexa AI assistant, which can turn on and off lights with voice commands. However, controlling the voice-activated robot with a higher degree of freedom presents several challenges, including the ambiguity in describing the scene and the potential for sentences to have multiple interpretations.
@@
Existing multimodal approaches have yet to meet expectations for general use, particularly for elderly users or patients who struggle with complex sign or command languages. A major challenge is the ineffective fusion of different information sources. Recent works on Visual Language Models (VLMs) [13] have been proposed to address this issue in HRI. However, these models are often application-specific or tailored to specific environments, lack comprehensive benchmarks against traditional methods, and require substantial GPU resources for processing. In contrast, large language models (LLMs) offer strong reasoning and emergent capabilities while maintaining reasonable computational requirements, making them promising candidates for effectively fusing and processing multimodal information.
Most research in robot action generation involves predefined domain searches, unconstrained exploration, behavior trees, and Bayesian inference. In one study, the authors integrated LLMs into robot action generation, defining a pipeline that converts human intentions into robotic action sequences using prompts and task-relevant APIs [14]. This approach provides a more intuitive and convenient method for user interaction and robot control. However, LLMs cannot directly obtain information from sensors, and Visual Language Models (VLMs) often require significant computational resources [13]. Furthermore, since LLMs are prone to hallucination, combining LLM with VLM increases the likelihood of errors and mistakes, often necessitating multiple trials for successful execution.
@@ VI-A Baseline Selections
In our work, we compare our approach with other open-source SOTA unimodal HRI methods [2, 6] and the multimodal method [13]. The baselines were selected primarily on the basis of similarity in action sequences and intent. The two unimodal benchmarks are gesture-based [2] and NLP-based [6], respectively. The multimodal approach is based on VLM [13], where target objects are selected through dialogue. The gesture-based HRI method [2], which utilizes the Leap Motion sensor, captures hand structure at specific localizations within a narrow field of view. The NLP and VLM approach [6, 13] demonstrated the use of language commands to direct the actions of a robot with a speech-to-text pipeline. NLP-based methods employ word embeddings, attention mechanisms, and probabilistic reasoning to recognize objects described in natural language. In contrast, VLM-based methods utilize the CLIP model to resolve ambiguities in object selection and assist users in forming clearer expressions. However, each baseline method has certain limitations and received numerous complaints from participants:
@@ VI-D Comparison of Robustness
As illustrated in 9(b), the system effectively captures human intention when the light is above 1 Lux. These findings confirm the adaptability of the system for most typical lighting conditions. However, in extreme low-light scenarios (e.g., At 1 lux or complete darkness), vision-based methods fail unless thermal imaging is used. While thermal imaging could address this issue, its high cost makes it impractical for elderly care applications.
@@ VII LIMITATION AND FUTURE WORKS
Our method enhances object detection using YOLO-World, a robust open-vocabulary model. However, its bias toward common knowledge limits accuracy for medical items. A potential solution is a database allowing nursing staff to update the system via verbal identification and online adaptation. Future work will explore this approach.

### v3

# Natural Multimodal Fusion-Based Human–Robot Interaction: Application With Voice and Deictic Posture via Large Language Model

@@ Abstract
publicationid: pubid: This work has been accepted for publication in IEEE Robotics and Automation Magazine (RAM) © 2025 IEEE.
Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses,
including reprinting/redistribution, creating new works, or reuse of any copyrighted components of this work in other media.
@@ I INTRODUCTION
With an aging population, the demand for efficient caregiving solutions is growing, yet labor costs continue to increase. Robotics manipulators [1] have shown promise in alleviating these challenges by automating tasks [2] that traditionally require human intervention [3, 4]. However, one of the most critical and unresolved issues is ensuring effective Human-Robot Interaction (HRI) [5] for elderly users. For HRI systems to truly benefit this population, they must offer interaction methods that are intuitive and natural, resembling everyday human communication. Current systems often rely on humans to memorize complex language syntax or master complex hand gestures [6, 7], which are impractical for the elderly. This highlights the urgent need for a simpler [8] yet highly effective method that allows robots to understand and execute commands from elderly users with ease and reliability.
In the past year, large language models (LLMs) [9, 10] have emerged as promising tools for HRI [11]. Their advanced reasoning and language capabilities make them promising for improving communication between humans and robots. However, directly applying LLMs to HRI presents several challenges. First, LLMs often require users to input detailed and structured text commands, which can be tedious and difficult to understand. Second, without integrated sensing capabilities, LLMs struggle to comprehend the environmental context [12, 13] or specific actions, limiting their effectiveness in real-world applications. Finally, LLMs are prone to hallucinations, generating inaccurate or unsafe responses, which can lead to harmful outcomes when used in control systems without close monitoring. These challenges highlight the need for a more robust integration of LLMs into HRI systems.
@@ II RELATED WORK
As perception and navigation solutions advance [14], robots are increasingly integrated into daily life. However, robots designed for complex tasks, such as surgical [15] and rehabilitation applications [16], face stringent safety requirements and must be validated for many years before widespread use in real-world scenarios. In contrast, service robots, particularly for elderly and patient care, demonstrate greater immediate potential. However, current service robot systems often rely on single-modal inputs, such as pure hand gestures [6], voice commands [17], or combinations of body language, such as hand gestures and body postures [18]. Hand gesture-based HRI methods [6], for example, frequently use the Leap Motion sensor to detect hand movements and translate them into commands [19]. However, these sensors have a limited field of view, restricting their applicability to broader environments. To address this, some researchers [18] have introduced Kinect cameras to extend the range of the sensor and detect both human posture and gestures. Nonetheless, for elderly users, accurately memorizing and executing the required gestures for commands remains a significant challenge, limiting the effectiveness of these approaches. Voice commands have always been a natural candidate for HRI problems. People often employ various approaches, such as LSTM, hidden Markov models, or an attention-based encoder-decoder network [20], to parse voices into actionable commands. One good example is the Amazon Alexa AI assistant, which can turn on and off lights with voice commands. However, controlling the voice-activated robot with a higher degree of freedom presents several challenges, including the ambiguity in describing the scene [21] and the potential for sentences to have multiple interpretations.
@@
Existing multimodal approaches have yet to meet expectations for general use, particularly for elderly users or patients who struggle with complex sign or command languages. A major challenge is the ineffective fusion of different information sources. Recent works on Visual Language Models (VLMs) [26] have been proposed to address this issue in HRI. However, these models are often application-specific or tailored to specific environments, lack comprehensive benchmarks against traditional methods, and require substantial GPU resources for processing. In contrast, large language models (LLMs) offer strong reasoning and emergent capabilities while maintaining reasonable computational requirements, making them promising candidates for effectively fusing and processing multimodal information [27].
Most research in robot action generation involves predefined domain searches, unconstrained exploration [28], behavior trees, and Bayesian inference. In one study, the authors integrated LLMs into robot action generation, defining a pipeline that converts human intentions into robotic action sequences using prompts and task-relevant APIs [29]. This approach provides a more intuitive and convenient method for user interaction and robot control [30]. However, LLMs cannot directly obtain information from sensors [31], and Visual Language Models (VLMs) often require significant computational resources [26]. Furthermore, since LLMs are prone to hallucination, combining LLM with VLM increases the likelihood of errors and mistakes, often necessitating multiple trials for successful execution.
@@ VI-A Baseline Selections
In our work, we compare our approach with other open-source SOTA unimodal HRI methods [6, 17] and the multimodal method [26]. The baselines were selected primarily on the basis of similarity in action sequences and intent. The two unimodal benchmarks are gesture-based [6] and NLP-based [17], respectively. The multimodal approach is based on VLM [26], where target objects [38] are selected through dialogue. The gesture-based HRI method [6], which utilizes the Leap Motion sensor, captures hand structure at specific localizations [39, 40, 41, 42] within a narrow field of view. The NLP and VLM approach [17, 26] demonstrated the use of language commands to direct the actions of a robot with a speech-to-text pipeline. NLP-based methods employ word embeddings, attention mechanisms, and probabilistic reasoning to recognize objects described in natural language. In contrast, VLM-based methods utilize the CLIP model to resolve ambiguities in object selection and assist users in forming clearer expressions. However, each baseline method has certain limitations and received numerous complaints from participants:
@@ VI-D Comparison of Robustness
As illustrated in 9(b), the system effectively captures human intention when the light is above 1 Lux. These findings confirm the adaptability of the system for most typical lighting conditions. However, in extreme low-light scenarios (e.g., At 1 lux or complete darkness), vision-based methods [43] fail unless thermal imaging is used. While thermal imaging could address this issue, its high cost makes it impractical for elderly care applications.
@@ VII LIMITATION AND FUTURE WORKS
Our method enhances object detection using YOLO-World, a robust open-vocabulary model. However, its bias toward common knowledge limits accuracy for medical items [44]. A potential solution is a database allowing nursing staff to update the system via verbal identification and online adaptation. Future work will explore this approach [45].